OpenStudy (anonymous):

can anyone pls answer this? solve and simplify the derivative of cot^2(y+2x)+ln[csc(y+2x)] = cos pi

OpenStudy (anonymous):

@lgbasallote i think your smart enough to solve this can you solve this for me?

OpenStudy (lgbasallote):

so you're supposed to take the derivative of that?

OpenStudy (anonymous):

yes @lgbasallote

OpenStudy (anonymous):

\[\frac{\cos^2(y+2x)}{\sin^2(y+2x)}+\ln(\frac{1}{\sin(y+2x)})=\cos(\pi)=-1\]

OpenStudy (anonymous):

First- use chain rule and differentiate wrt (y+2x). -For the cotangent use the product rule for the denominator and numerator -For the natural log use the chain rule a couple of times.

OpenStudy (anonymous):

Do you have any idea how to start on this?

OpenStudy (anonymous):

@happy87 ?

OpenStudy (anonymous):

yes @henpen i'm sorry for the late reply

OpenStudy (anonymous):

can you show me the whole solution @henpen ? thanks that would be very helpful

OpenStudy (anonymous):

Differenitate both sides wrt t: -1 turns to 0.

OpenStudy (anonymous):

Addition rule or whatever it's called- the two terms on each side of the + (or -) differentiate independently.

OpenStudy (anonymous):

Let's work on the right one first.

OpenStudy (anonymous):

\[\frac{dln\frac{1}{\sin(y+2x)}}{dx}= \frac{dln\frac{1}{\sin(y+2x)}}{d\frac{1}{\sin(y+2x)}} \frac{ d\frac{1}{\sin(y+2x)} }{dsin(y+2x)}\frac{dsin(y+2x)}{d(y+2x)}\frac{d(y+2x)}{dx}\]

OpenStudy (anonymous):

So, for example, treat the first term in the multiplication as